A FASCINATING THESIS

A thesis submitted for the degree of Doctor of Philosophy

By Fred Nerk B.Sc.

School of Physics, University College, The University of New South Wales, Australian Defence Force Academy.

November 1999

Abstract

This thesis tells you all you need to know about...

Acknowledgements

I would like to thank...

Contents

Abstract i Acknowledgements iii References 1

List of Figures

vii

References

1